# OpenSocInt: A Multi-modal Training Environment for Human-Aware Social Navigation

### v1

@@ 2.1 The Simulator Module
In order to compute the updated state, the simulator needs to compute the next position of the agent (given the action) and the next position of the dynamic obstacles (persons). We hypothesise that each person has a goal, and the next position of each person will be computed using a weighted sum of a social force model [pedica2008social] and a goal force model. While for the later, the force is oriented toward the goal and proportional to the distance (with maximum unitary value), the later is based on a repulsion force to avoid humans to be too close to each other, see Figure 3. This repulsion force does not take the agent into account as preliminary experiments showed that the agent will learn to exploit this repulsion to push humans away and go directly to its goal.
@@ 2.3 The Agent Module
The multi-modal agent architecture inputting the various modalities is shown in Figure 4. A latent space (blue box) is learn from one or more modalities plus the agent’s goal. The local egocentric occupancy grid (LEOG) is processed by a few convolutional layers (in pink), then flattened and transformed via one or more fully connected layers (in green). The LIDAR (RayCast) and the other two modalities are transformed via one or more fully connected layers. In order to facilitate learning from very low-dimensional inputs (e.g. the relative position of the goal and closest obstacle) a radial basis function [montazer2018radial] layer (in orange) is used to increase the representation’s dimensionality before the fully connected layers. The one or multiple modalities are concatenated into the latent representation, that is then used by the various reinforcement learning agents implemented.
All the implemented agents exploit the same latent representation (blue in the image) and cover the following deep RL paradigms: twin-delayed deep deterministic policy gradient (TD3) [dankwa2019twin], deep deterministic policy gradient (DDPG) [lillicrap2015continuous], advantage actor-critic (A2C) [kumar2023sample] and soft-actor critic (SAC) [haarnoja2018soft].
@@ 3.2 Metrics
We report the evolution of three metrics along the training. First, the percentage of episodes that end with the agent reaching the goal. Second, the percentage of episodes that end with the agent colliding eiguillaumesire@gmail.comther with a static or with a dynamic obstacle. Third, the percentage of episodes that run until the maximum number of steps without reaching the goal or colliding, see the previous section. Importantly, these metrics are computed in a set of episodes that does not belong to the training set, and is therefore never part of the replay buffer. More precisely, every 50 training episodes, the model is evaluated on 20 test episodes, until reaching 700 training episodes (or 280 test episodes). This is done 5 times per configuration. The mean and standard deviation of each metric over a sliding window of 10 test episodes is the reported.

### v2

@@ 2.1 The Simulator Module
In order to compute the updated state, the simulator needs to compute the next position of the agent (given the action) and the next position of the dynamic obstacles (persons). We hypothesise that each person has a goal, and the next position of each person will be computed using a weighted sum of a social force model [6] and a goal force model. While for the later, the force is oriented toward the goal and proportional to the distance (with maximum unitary value), the later is based on a repulsion force to avoid humans to be too close to each other, see Figure 3. This repulsion force does not take the agent into account as preliminary experiments showed that the agent will learn to exploit this repulsion to push humans away and go directly to its goal.
@@ 2.3 The Agent Module
The multi-modal agent architecture inputting the various modalities is shown in Figure 4. A latent space (blue box) is learn from one or more modalities plus the agent’s goal. The local egocentric occupancy grid (LEOG) is processed by a few convolutional layers (in pink), then flattened and transformed via one or more fully connected layers (in green). The LIDAR (RayCast) and the other two modalities are transformed via one or more fully connected layers. In order to facilitate learning from very low-dimensional inputs (e.g. the relative position of the goal and closest obstacle) a radial basis function [5] layer (in orange) is used to increase the representation’s dimensionality before the fully connected layers. The one or multiple modalities are concatenated into the latent representation, that is then used by the various reinforcement learning agents implemented.
All the implemented agents exploit the same latent representation (blue in the image) and cover the following deep RL paradigms: twin-delayed deep deterministic policy gradient (TD3) [1], deep deterministic policy gradient (DDPG) [4], advantage actor-critic (A2C) [3] and soft-actor critic (SAC) [2].
@@ 3.2 Metrics
We report the evolution of three metrics along the training. First, the percentage of episodes that end with the agent reaching the goal. Second, the percentage of episodes that end with the agent colliding either with a static or with a dynamic obstacle. Third, the percentage of episodes that run until the maximum number of steps without reaching the goal or colliding, see the previous section. Importantly, these metrics are computed in a set of episodes that does not belong to the training set, and is therefore never part of the replay buffer. More precisely, every 50 training episodes, the model is evaluated on 20 test episodes, until reaching 700 training episodes (or 280 test episodes). This is done 5 times per configuration. The mean and standard deviation of each metric over a sliding window of 10 test episodes is the reported.